Jane has 116 more marbles than Mary has. If their mother gives 18 marbles to each of them, Jane will have 5 times as many marbles as Mary. How many marbles does Mary have before her mother gives them any more?
Answer by ankor@dixie-net.com(12704)

Jane has 116 more marbles than Mary has. If their mother gives 18 marbles to each of them, Jane will have 5 times as many marbles as Mary. How many marbles does Mary have before her mother gives them any more?
:
Let x = number of marbles Mary has. [that's what they are asking for]
:
Write an algebra expression for each phrase:
:
"Jane has 116 more marbles than Mary"
Jane's marbles = [x + 116]
:
"Their mother give 18 marbles to each"
Jane:
[x+116] + 18
[x + 134]
:
Mary:
[x + 18]
:
:
"Jane will have 5 times as many marbles as Mary."
x + 134 = 5[x + 18]
x + 134 = 5x + 90
134 - 90 = 5x - x
44 = 4x
x = 44/4
x = 11 marbles for Mary to start with
:
:
Check: Jane: 116 + 11 = 127; Mom adds 18 for 145 total marbles
Mary: 11 + 18 that Mom gave her for 29 total.
5 * 29 = 145 which is Janes total